OpenStudy (anonymous):

Given the linear equation y= -2/7x+3, find the y-coordinates of the points (-7, ), (0, ), and (7, ). Please show all of your work.

OpenStudy (anonymous):

plug in the x coordiante and you will get the corresponding y coordinate

OpenStudy (anonymous):

i did that and still not gettin the correct answer

OpenStudy (anonymous):

(7, ) y=(-2/7(7))+(3) y=(-2/49)+(3) y=145/49 (7,145/49)

OpenStudy (anonymous):

Thats the first one, can you the other one from that?

OpenStudy (anonymous):

im trying right now

OpenStudy (anonymous):

the first and the lanes ones are suppose to be the same?

OpenStudy (anonymous):

no actually the one i did was tha last one. The other tow should be different